1/8(5Y+64)=1/4(20+2Y) Solve the equation...If possible
Answer by rfer(16322)

1/8(5y+64)=1/4(20+2y)
0.625y+8=5+0.5y
0.125y=-3
y=24
------------------
5/8y+64/8=20/4+2/4y
-1/8y=24/8
y=24/8*-8/1
y=-24